# Law of Large Numbers for continuous $N$-particle ensembles at fixed temperature

### v1

@@ 4.2 Proof of the “if” part of the main theorem
Let us describe the proof strategy of (23). In Step 1, we discuss the combinatorial behavior that different parts of the operators have. In Step 2, we decompose the action as a sequence of simple transformations on polynomials, and (23) is a combination of such sequences. In Step 3 and 4, we identify the collection of sequences that gives a nontrivial weight contribution in the limit. Finally, in Step 5, we explicitly calculate the asymptotic of Eqn. (23).
Step 1. Let $h(\vec{x})$ be a polynomial of $x_{1},\dots,x_{N}$, let $\kappa$ be any partition, and recall that $p_{\kappa}(\vec{x})=\prod_{i=1}^{\ell(\kappa)}p_{\kappa_{i}}(\vec{x})$. Straightforward calculations yield the following equalities:
$\partial_{i}\left(h(\vec{x})\,p_{\kappa}(\vec{x})\,e^{F_{N,\theta}(\vec{x})}\right)=$ $\ \left[\partial_{i}h(\vec{x})+\sum_{s=1}^{\ell(\kappa)}{\kappa_{s}x_{i}^{\kappa_{s}-1}\prod_{j\neq s}{p_{\kappa_{j}}(\vec{x})}}\right]\cdot e^{F_{N,\theta}(\vec{x})}$
$+\bigg(\sum_{d=1}^{\infty}{da_{(d)}^{N}x_{i}^{d-1}}\bigg)h(\vec{x})\,p_{\kappa}(\vec{x})\ e^{F_{N,\theta}(\vec{x})}$
$+\bigg(\sum_{\lambda\colon\ell(\lambda)\geq 2}{a_{\lambda}^{N}\sum_{j=1}^{\ell(\lambda)}{\lambda_{j}x_{i}^{\lambda_{j}-1}\prod_{k\neq j}{p_{\lambda_{k}}(\vec{x})}}}\bigg)h(\vec{x})\,p_{\kappa}(\vec{x})e^{F_{N,\theta}(\vec{x})},$
$\frac{1-\sigma_{ij}}{x_{i}-x_{j}}\left(h(\vec{x})\,p_{\kappa}(\vec{x})\,e^{F_{N,\theta}(\vec{x})}\right)=\frac{1-\sigma_{ij}}{x_{i}-x_{j}}\big(h(\vec{x})\big)p_{\kappa}(\vec{x})\,e^{F_{N,\theta}(\vec{x})},$
$\frac{1-\sigma_{ij}}{x_{i}-x_{j}}\Big(x_{i}^{m_{i}}x_{j}^{m_{j}}\Big)=\begin{cases}-\sum_{a=m_{i}}^{m_{j}-1}{x_{i}^{a}x_{j}^{m_{i}+m_{j}-1-a}},&\text{if }m_{i}<m_{j},\\ \\ \sum_{a=m_{j}}^{m_{i}-1}{x_{i}^{a}x_{j}^{m_{i}+m_{j}-1-a}},&\text{if }m_{i}\geq m_{j}.\end{cases}$
Step 2. We introduce the following linear operators on polynomials, uniquely defined by their actions on monomials $x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}$, for any $m_{1},...,m_{N}\in\mathbb{Z}_{\geq 0}$.
  • For $1\leq i\leq s$, $1\leq d\leq|\mu|$,
    $U^{(d)}_{i}(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}):=da^{N}_{(d)}x_{i}^{d-1}\cdot x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}.$
  • For $1\leq i\leq s$, and any partition $\lambda$ such that $\ell(\lambda)\geq 2$ and $|\lambda|\leq|\mu|$,
    $U_{i}^{(\lambda)}(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}):=a_{\lambda}^{N}\sum_{j=1}^{\ell(\lambda)}{\lambda_{j}x_{i}^{\lambda_{j}-1}\prod_{k\neq j}{p_{\lambda_{k}}(\vec{x})}}\cdot x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}.$
  • For $1\leq i\leq s$, $1\leq j\leq N$, $j\neq i$,
    $d_{i,j}(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}):=\begin{cases}x_{1}^{m_{1}}\cdots x_{i-1}^{m_{i-1}}x_{i}^{m_{i}-1}x_{i+1}^{m_{i+1}}\cdots x_{N}^{m_{N}},&m_{i}>m_{j};\\ 0,&m_{i}\leq m_{j}.\end{cases}$
  • For $1\leq i\leq s$, $1\leq j\leq N$, $j\neq i$,
    $\tilde{d}_{ij}(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}):=\begin{cases}-x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot\sum_{a=1}^{m_{j}-m_{i}}{x_{i}^{a-1}x_{j}^{-a}},&\text{if }m_{i}<m_{j},\\ x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot\sum_{a=1}^{m_{i}-m_{j}-1}{x_{i}^{-a-1}x_{j}^{a}},&\text{if }m_{i}\geq m_{j}.\end{cases}$
  • For $1\leq i\leq s$,
    $\partial_{i}(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}})=m_{i}\cdot d_{i}(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}),$
    where
    $d_{i}(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}):=\mathbf{1}_{\{m_{i}\geq 1\}}\cdot x_{1}^{m_{1}}\cdots x_{i-1}^{m_{i-1}}x_{i}^{m_{i}-1}x_{i+1}^{m_{i+1}}\cdots x_{N}^{m_{N}}.$
Then, according to the formulas from Step 1, we have
$\begin{split}&N^{-|\mu|}\Big[\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{2}^{\mu_{2}}\mathcal{D}_{1}^{\mu_{1}}\Big]e^{F_{N,\theta}(\vec{x})}\,\Big|_{\vec{x}=(0^{N})}\\ =&N^{-|\mu|}\left(\sum_{d}U^{(d)}_{s}+\sum_{\lambda}U^{(\lambda)}_{s}+\theta\sum_{j\neq s}d_{s,j}+\theta\sum_{j\neq s}\tilde{d}_{s,j}+\partial_{s}\right)^{\mu_{s}}\cdots\\ &\left(\sum_{d}U^{(d)}_{1}+\sum_{\lambda}U^{(\lambda)}_{1}+\theta\sum_{j\neq 1}d_{1,j}+\theta\sum_{j\neq 1}\tilde{d}_{1,j}+\partial_{1}\right)^{\mu_{1}}(1)\\ =&N^{-|\mu|}\sum_{T(l):1\leq l\leq|\mu|}\prod_{l=1}^{|\mu|}T(l)(1),\end{split}$ (24)
where the summation in the last line is over the finitely many choices of the operators $T(l)$. More precisely, $T(l)$ is chosen among $U^{(d)}_{i}$’s, $U^{(\lambda)}_{i}$’s, $\theta\cdot d_{i,j}$’s, $\theta\cdot\tilde{d}_{i,j}$’s and $\partial_{i}$ for $\sum_{j=1}^{i-1}\mu_{j}+1\leq l\leq\sum_{j=1}^{i}\mu_{j}$. We call the above five types of operators are of Types I to V, respectively, and denote the set of Type $i$ operators by $J_{i}\ (i=I,\dots,V)$. For a given sequence $\prod_{l=1}^{|\mu|}T(l)$, denote the number of $T(l)\in J_{i}$ by $M_{i}$. Note that the degree of the polynomial increases or is unchanged under the action of Type I or II, and decreases by 1 under Type III to V actions.
Denote $\tilde{T}:=(\tilde{T}(|\mu|),\dots,\tilde{T}(1))$, and let $C(n)$ denote the constant term of $\prod_{l=1}^{n}T(l)(1)$ (we omit the dependence on $T(l)$’s in the notation). We will treat $n$ as the occurring time of the action $T(n)$. Given $\vec{T}$, we say that $T(n)$ acts as 0, if $\prod_{l=1}^{n-1}T(l)(1)\neq 0$, $\prod_{l=1}^{n}T(l)(1)=0$.
Claim 1. For each $1\leq l\leq|\mu|$, there is a constant $K(\mu)$ uniform in $N$, such that $|J_{I}|+|J_{II}|+|J_{V}|$, and the total degree of $\prod_{l=1}^{n}T(l)(1),\ 1\leq n\leq|\mu|$, is bounded by $K(\mu)$. Moreover, $|J_{III}|\leq sN$, $|J_{IV}|\leq sN$.
Proof of the claim. Since we assume that $F_{N,\theta}(\vec{x})$ is a polynomial of degree $|\mu|$, the total number of operators $U^{(d)}_{i}$’s and $U^{(\lambda)}_{i}$’s is bounded. The total degree is bounded since there are finitely many iterations of finitely many choices of $T(l)\in J_{I}\cup J_{II}$, such that each choice increases the degree by at most $|\mu|-1$. The second statement is clear from definition.∎
Claim 2. If $T(l_{1})=\theta\tilde{d}_{i_{1},j}$ for some $1\leq l_{1}\leq|\mu|$, $1\leq i_{1}\leq s$, $s+1\leq j\leq N$, then $C(|\mu|)\neq 0$ only if for some $l_{2}>n_{1}$, $T(l_{2})=d_{i_{2},j}$.
Proof of the claim. Without loss of generality, assume that $n_{1}$ is the first time that $T(l)=\tilde{d}_{i,j}$ for this given $j$. By definition of ${d}_{i,j}$, either $\prod_{l=1}^{n_{1}}T(l)(1)=0$, or $deg(x_{j})>0$ in every monomial in $\prod_{l=1}^{n_{1}}T(l)(1)$. In either case, again by definition, the action $\prod_{l=n_{1}+1}^{|\mu|}T(l)$ where $T(l)\notin\{\tilde{d}_{1,j},\tilde{d}_{s,j}\}$ for all $l$’s can only send $\prod_{l=1}^{n_{1}}T(l)(1)$ to 0.∎
Step 3. We consider two sub-cases of the sequence. Firstly, assume that there are no Type II operators in the sequence.
Claim 3: We have
$\lim_{N\rightarrow\infty}N^{-|\mu|}\sum_{\vec{T}:\,M_{II}=0,M_{IV}+M_{V}\geq 1}\left|C(|\mu|)\right|=0.$
Proof of the claim: We divide $\vec{T}$’s into finitely many collections where each collection shares the same data $(M_{I},...,M_{V})$. Then it suffices to verify the claim for each group. By Claim 1, the total number of sequences $\prod_{l=1}^{|\mu|}T(l)$ with $M_{i}$ operators in Type $i$ is at most
$K^{M_{I}+M_{V}}(sN)^{M_{III}+M_{IV}}.$
By Claim 2, if $T(l_{1})\in J_{IV}$ for some $l_{1}<|\mu|$, then there are at most $s$ choices of $T(l_{2})$ in one of the $l_{1}<l_{2}\leq|\mu|$, so that $C(|\mu|)\neq 0$. We also observe from the definition that if $T(|\mu|)\in J_{IV}$, then $C(|\mu|)=0$. Therefore, the total number of sequences $\prod_{l=1}^{|\mu|}T(l)$ with $M_{i}$ operators in Type $i$, such that $C(|\mu|)\neq 0$, is bounded by
$K^{M_{I}+M_{V}}(sN)^{M_{III}+M_{IV}-1}\cdot s|\mu|.$ (25)
In addition, each $T(l)\in J_{I}\cup J_{III}\cup J_{V}$ turns a monomial into another monomial. Regarding $T(l)\in J_{IV}$, note that the number of monomials in the output is bounded by the total degree of the input, which is bounded by $K$ by Claim 1. Therefore, we treat $\{N^{-|\mu|}\prod_{l=1}^{n}T(l)(1)\}_{n=1}^{|\mu|}$ as sum of sequences of monomials with no more than $K^{|\mu|}$ summands.
For each such sequence of monomials, its coefficient is bounded in absolute value by
$N^{-|\mu|}\prod_{d=1}^{|\mu|}(da^{N}_{d})^{\#\{1\leq l\leq s\colon T(l)=U^{(d)}_{i}\,\text{for\ some\ }i\}}\cdot\theta^{M_{III}+M_{IV}}K^{M_{V}}.$ (26)
Recall that $|\mu|=M_{I}+\cdots+M_{V}.$ Multiplying Eqn. (25) and (26), and by assumption (a) in Definition 3.2, we have that
$\begin{split}&N^{-|\mu|}\sum_{M_{II}=0,M_{IV}+M_{V}\geq 1}\left|C(|\mu|)\right|\\ \leq&K^{M_{I}+M_{V}}(sN)^{M_{II}+\max(M_{IV}-1,0)}\cdot s|\mu|N^{-|\mu|}\prod_{d=1}^{|\mu|}(da^{N}_{(d)})^{\#\{1\leq l\leq s\colon T(l)=U^{(d)}_{i}\ \text{for\ some\ }i\}}\cdot\theta^{M_{III}+M_{IV}}K^{M_{V}}\\ \lesssim&N^{M_{I}+M_{II}+\max(M_{IV}-1,0)-|\mu|}=N^{-M_{V}-\mathbf{1}_{\{M_{IV}\geq 1\}}}.\end{split}$ (27)
Step 4. Secondly, we show that the contribution of sequences $\vec{T}$ with Type II operators also vanishes. To achieve this, we introduce another collection of linear operators that acts on the polynomial of the form $const\times x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot p_{\kappa}(\vec{x})$, where $m_{1},\dots m_{N}\in Z_{\geq 0}$, $\kappa$ is a partition.
  • For $1\leq i\leq s$, $1\leq d\leq|\mu|$,
    $U^{(d)}_{i}(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot p_{\kappa}(\vec{x})):=da^{N}_{(d)}x_{i}^{d-1}\cdot x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot p_{\kappa}(\vec{x}).$
  • For $1\leq i\leq s$, and any partition $\lambda$ such that $\ell(\lambda)\geq 2$ and $|\lambda|\leq|\mu|$,
    $U_{i}^{(\lambda)}(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot p_{\kappa}(\vec{x}))=a_{\lambda}^{N}\sum_{j=1}^{\ell(\lambda)}{\lambda_{j}x_{i}^{\lambda_{j}-1}\prod_{j\neq i}{p_{\lambda_{j}}(\vec{x})}}\cdot x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot p_{\kappa}(\vec{x}).$
  • For $1\leq i\leq s$,
    $T_{\sigma,i}\left(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot p_{\kappa}(\vec{x})\right)=\sum_{j\neq i}(d_{i,j}+\tilde{d}_{i,j})\left(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\right)\cdot p_{\kappa}(\vec{x}).$
  • For $1\leq i\leq s$,
    $\partial_{i,1}(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot p_{\kappa}(\vec{x})):=m_{i}\cdot d_{i}(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}})\cdot p_{\kappa}(\vec{x}).$
  • For $1\leq i\leq s$,
    $\partial_{i,2}(x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot p_{\kappa}(\vec{x})):=x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot\left(\sum_{s=1}^{\ell(\kappa)}{\kappa_{s}x_{i}^{\kappa_{s}-1}\prod_{j\neq s}{p_{\kappa_{j}}(\vec{x})}}\right).$
Then, according to the discussion in Step 1, we have
$\begin{split}&N^{-|\mu|}\Big[\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{2}^{\mu_{2}}\mathcal{D}_{1}^{\mu_{1}}\Big]e^{F_{N,\theta}(\vec{x})}\,\Big|_{\vec{x}=(0^{N})}\\ =&N^{-|\mu|}\left(\sum_{d}U^{(d)}_{s}+\sum_{\lambda}U^{(\lambda)}_{s}+\theta\,T_{\sigma,s}+\partial_{s,1}+\partial_{s,2}\right)^{\mu_{s}}\cdots\\ &\left(\sum_{d}U^{(d)}_{1}+\sum_{\lambda}U^{(\lambda)}_{1}+\theta\,T_{\sigma,1}+\partial_{1,1}+\partial_{1,2}\right)^{\mu_{1}}(1)\\ =&N^{-|\mu|}\sum_{\tilde{T}(l):1\leq l\leq|\mu|}\prod_{l=1}^{|\mu|}\tilde{T}(l)(1),\end{split}$ (28)
where the summation in the last line is over the finitely many choices of the operators $\tilde{T}(l)$. More precisely, $\tilde{T}(l)$ is chosen among $U^{(d)}_{i}$’s, $U^{(\lambda)}_{i}$’s, $\theta\cdot T_{\sigma,i}$, $\partial_{i,1}$ and $\partial_{i,2}$ for $\sum_{j=1}^{i-1}\mu_{j}+1\leq l\leq\sum_{j=1}^{i}\mu_{j}$. We call the above five types of operators are of Types A to E [^2] By definition, Type A, B, D are the same as Type I, II, V respectively, and Type C is the union of Type III, IV., respectively, and denote the set of Type $i$ operators by $J_{i}\ (i=A,\dots,E)$. For a given sequence $\prod_{l=1}^{|\mu|}\tilde{T}(l)$, denote the number of $\tilde{T}(l)\in J_{i}$ by $M_{i}$. Note that the degree of the polynomial increases or is unchanged under the action of Type A or B, and decreases by 1 under Type C to E actions.
Denote $\tilde{T}:=(\tilde{T}(|\mu|),\dots,\tilde{T}(1))$, and let $\tilde{C}(|\mu|)$ denote the constant term of $\prod_{l=1}^{|\mu|}\tilde{T}(l)(1)$ (we omit the dependence on $\tilde{T}(l)$’s in the notation).
Claim 4: We have
$\lim_{N\rightarrow\infty}N^{-|\mu|}\sum_{M_{B}\geq 1}\left|C(|\mu|)\right|=0.$
Proof of the the claim: We divide $\vec{\tilde{T}}$’s into finitely many collections where each collection shares the same data $(M_{A},...,M_{E})$. Then it suffices to verify the claim for each group. Suppose that for $1\leq l_{1},\dots l_{n}<|\mu|$, $\tilde{T}(l_{m})=U^{(\lambda(m))}_{i_{m}}\ (1\leq m\leq n)$ is the collection of $\tilde{T}(l)\in J_{B}$, then we claim that $\tilde{C}(|\mu|)=0$, unless $M_{E}\geq\sum_{m=1}^{n}l(\lambda(m))-n$. Indeed, each $U^{(\lambda(m))}_{i_{m}}$ creates (a finite [^3] The number of summands is again bounded by the uniform constant $K$. sum of) $l(\lambda(m))-1$ copies of power sums factor of the form $\prod_{j\neq s}p_{\lambda(m)_{j}}(\vec{x})$. To a nonzero constant after $|\mu|$ steps, one needs to eliminate of of these power sums. Since power sums are invariant under the action of $T_{\sigma,i}$, this can only be achieved by $\partial_{i,2}$’s.
More precisely, each copy of $\tilde{T}(l)\in J_{E}$ removes one copy of power sum in $x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot p_{\kappa}(\vec{x})$. On the other hand, $\tilde{T}(l)\in J_{A}\cup J_{C}\cup J_{D}$ acts only on $x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}$. By definition, one can see that Type A, B, D, E operators convert a monomial to at most $K$ monomials, and Type C operators provide at most $2sN$ monomials. Therefore, up to a constant independent of $N$, the total number of monomials in $N^{-|\mu|}\prod_{l=1}^{|\mu|}\tilde{T}(l)(1)$ is bounded by
$N^{M_{C}}.$ (29)
For each such sequence of polynomials of the form $const\times x_{1}^{m_{1}}\cdots x_{N}^{m_{N}}\cdot p_{\kappa}(\vec{x})$, its coefficient is bounded in absolute value by
[…]

### v2

@@ 4.2 Proof of the “if” part of the main theorem
Step 1. Let $\mu:=(\mu_{1},\dots,\mu_{s})$ and denote $|\mu|:=\mu_{1}+\dots+\mu_{s}$. Note that $\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{1}^{\mu_{1}}$ is a sum of $N^{|\mu|}$ terms $w^{s}_{\mu_{s}}\cdots w^{s}_{1}\cdots w^{1}_{\mu_{1}}\cdots w^{1}_{1}$, where each $w^{a}_{b}$ is of the form $\partial_{a}$, or $\frac{\theta}{x_{a}-x_{j}}(1-\sigma_{a,j})$, for some $j\neq a$. Before taking the limit, we need to compute $\Big[\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{1}^{\mu_{1}}\Big]e^{F_{N,\theta}(\vec{x})}$, which is the sum of all terms $\big(w^{s}_{\mu_{s}}\cdots w^{s}_{1}\cdots w^{1}_{\mu_{1}}\cdots w^{1}_{1}\big)e^{F_{N,\theta}(\vec{x})}$. If $w^{1}_{1}$ is of the form $\frac{\theta}{x_{1}-x_{j}}(1-\sigma_{1,j})$, then it kills $e^{F_{N,\theta}(\vec{x})}$, for being symmetric in all variables $x_{i}$, therefore we can ignore those terms and keep only the terms with $w^{1}_{1}=\partial_{1}$. Evidently, $\partial_{1}e^{F_{N,\theta}(\vec{x})}=\big(\partial_{1}F_{N,\theta}(\vec{x})\big)e^{F_{N,\theta}(\vec{x})}$. After that, we apply the operators $w^{1}_{2},w^{1}_{3},\dots,w^{s}_{\mu_{s}}$ (in that order) to expressions of the form $h(\vec{x})e^{F_{N,\theta}(\vec{x})}$, for some polynomial $h(\vec{x})$, and keep obtaining expressions of this form, as observed by the equalities:
$\partial_{i}\left(h(\vec{x})\,e^{F_{N,\theta}(\vec{x})}\right)=\partial_{i}h(\vec{x})e^{F_{N,\theta}(\vec{x})}+\left(\sum_{\nu\colon|\nu|\leq|\mu|}\ \sum_{j=1}^{\ell(\nu)}{\nu_{j}a_{\nu}^{N}x_{i}^{\nu_{j}-1}\prod_{k\neq j}{p_{\nu_{k}}(\vec{x})}}\right)h(\vec{x})\,e^{F_{N,\theta}(\vec{x})},$ (24)
$\theta\cdot\frac{1-\sigma_{a,j}}{x_{a}-x_{j}}\left(h(\vec{x})\,e^{F_{N,\theta}(\vec{x})}\right)=\ \theta\cdot\frac{1-\sigma_{a,j}}{x_{a}-x_{j}}\big(h(\vec{x})\big)\,e^{F_{N,\theta}(\vec{x})}.$
Then, to obtain the limit of $\Big[\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{1}^{\mu_{1}}\Big]e^{F_{N,\theta}(\vec{x})}$, we proceed as follows.
First, expand each term $\big(w^{s}_{\mu_{s}}\cdots w^{s}_{1}\cdots w^{1}_{\mu_{1}}\cdots w^{1}_{1}\big)e^{F_{N,\theta}(\vec{x})}$ (by the above rules) as a sum that can be finally be expressed as $H(\vec{x})e^{F_{N,\theta}(\vec{x})}$, for some polynomial $H(\vec{x})$. This will also be a polynomial in $N,\theta$, as well as $\{a_{(d)}^{N}:d\in\mathbb{Z}_{\geq 1}\}$ and $\{a_{\lambda}^{N}:\ell(\lambda)\geq 2\}$.
Second, we set $\vec{x}\mapsto(0^{N})$. Since $F_{N,\theta}(\vec{x})\big|_{\vec{x}=(0^{N})}=0$, then $H(\vec{x})e^{F_{N,\theta}(\vec{x})}\big|_{\vec{x}=(0^{N})}=H(0^{N})$ is the constant coefficient of $H(\vec{x})$, which is a polynomial in $N$, $\theta$, $\{a_{(d)}^{N}:d\in\mathbb{Z}_{\geq 1}\}$ and $\{a_{\lambda}^{N}:\ell(\lambda)\geq 2\}$.
Finally, we multiply by $N^{-|\mu|}$ and take the limit of the resulting expression as $N\to\infty$, using assumptions (16). The result will be a polynomial in $\theta$ and $\{\kappa_{d}:d\in\mathbb{Z}_{\geq 1}\}$. We need to prove that the final result is the right hand side of (23).
Step 2. Let $\nu^{(1)},\dots,\nu^{(k)}$ be partitions such that constant multiples of the product $a_{\nu^{(1)}}^{N}\cdots a_{\nu^{(k)}}^{N}$ appear in terms coming from some $\big(w^{s}_{\mu_{s}}\cdots w^{s}_{1}\cdots w^{1}_{\mu_{1}}\cdots w^{1}_{1}\big)e^{F_{N,\theta}(\vec{x})}$. We want to calculate the contribution of those terms after setting $\vec{x}\mapsto(0^{N})$, multiplying by $N^{-|\mu|}$, and taking the limit $N\to\infty$. By the rule (24), we see that $a_{\nu^{(r)}}^{N}$ appears for the first time after an application of some $\partial_{i}$, as part of the expression
$\nu^{(r)}_{j}a_{\nu^{(r)}}^{N}x_{i}^{\nu^{(r)}_{j}-1}\prod_{k\neq j}{p_{\nu^{(r)}_{k}}(\vec{x})}\cdot e^{F_{N,\theta}(\vec{x})}.$ (25)
Since we want the product $a_{\nu^{(1)}}^{N}\cdots a_{\nu^{(k)}}^{N}$, we need $k$ partial derivatives among the $w^{a}_{b}$ in order to get the $k$ factors $a^{N}_{\nu^{(r)}}$, $1\leq r\leq k$. Then we have another $|\mu|-k$ operators of the form $\partial_{a}$ or $\frac{\theta}{x_{a}-x_{j}}(1-\sigma_{a,j})$ that must act on expressions $h(\vec{x})e^{F_{N,\theta}(\vec{x})}$ (but $\partial_{a}$ would act on $h(\vec{x})$ and not on $e^{F_{N,\theta}(\vec{x})}$) and each of them decreases the degree of $h(\vec{x})$ by $1$. Then we set $\vec{x}=(0^{N})$, so we want the expressions $h(\vec{x})e^{F_{N,\theta}(\vec{x})}$ with $h(\vec{x})$ being a constant at the end.
We note that if $g(\vec{x})$ is a symmetric function in the variables $x_{i}$, then $\frac{\theta}{x_{a}-x_{j}}(1-\sigma_{a,j})\big(g(\vec{x})h(\vec{x})\big)=g(\vec{x})\cdot\frac{\theta}{x_{a}-x_{j}}(1-\sigma_{a,j})\big(h(\vec{x})\big)$. This implies that if $\frac{\theta}{x_{a}-x_{j}}(1-\sigma_{a,j})$ is applied to (25), then each of the factors $p_{\nu_{k}^{(r)}}(\vec{x})$, where $1\leq k\leq\ell(\nu^{(r)})$ and $k\neq j$, is preserved. But, in the end, these factors cannot be present, or otherwise setting $\vec{x}\mapsto(0^{N})$ would make its contribution vanishing. Thus, we need $\ell(\nu^{(r)})-1$ partial derivatives to act on (25). Hence, for $a_{\nu^{(1)}}^{N}\cdots a_{\nu^{(k)}}^{N}$ to appear inside some $\big(w^{s}_{\mu_{s}}\cdots w^{s}_{1}\cdots w^{1}_{\mu_{1}}\cdots w^{1}_{1}\big)e^{F_{N,\theta}(\vec{x})}$, one must have at least $k+\sum_{r=1}^{k}{(\ell(\nu^{(r)})-1)}=\sum_{r=1}^{k}{\ell(\nu^{(r)})}$ partial derivatives among the $w^{a}_{b}$, while the other $|\mu|-\sum_{r=1}^{k}{\ell(\nu^{(r)})}$ operators can be any of the $N$ operators $\partial_{a}$ or $\frac{\theta}{x_{a}-x_{j}}(1-\sigma_{a,j})$, for some $1\leq j\leq N$, $j\neq a$.
As a result of the previous analysis, the coefficient of $a_{\nu^{(1)}}^{N}\cdots a_{\nu^{(k)}}^{N}$ in $\Big[\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{1}^{\mu_{1}}\Big]e^{F_{N,\theta}(\vec{x})}\,\big|_{\vec{x}=(0^{N})}$ is of order $O\big(N^{|\mu|-\sum_{r=1}^{k}{\ell(\nu^{(r)})}}\big)$, so the coefficient of $a_{\nu^{(1)}}^{N}\cdots a_{\nu^{(k)}}^{N}$ in $N^{-|\mu|}\,\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{1}^{\mu_{1}}e^{F_{N,\theta}(\vec{x})}\,\big|_{\vec{x}=(0^{N})}$ is of order $O\big(N^{-\sum_{r=1}^{k}{\ell(\nu^{(r)})}}\big)$. By the assumptions (16), the limit $\lim_{N\to\infty}{N^{-\ell(\nu^{(r)})}a^{N}_{\nu^{(r)}}}$ is finite for all partitions $\nu^{(r)}$, and it is zero when $\ell(\nu^{(r)})\geq 2$. The conclusion is that the limit in the left hand side of (23) exists and is finite and that we can ignore terms that contain some $a^{N}_{\lambda}$, with $\ell(\lambda)\geq 2$, in the prelimit expression $\Big[\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{1}^{\mu_{1}}\Big]e^{F_{N,\theta}(\vec{x})}$, since their final contribution is zero.
Step 3. In Step 2, we showed that the prelimit expression in the left hand size of (23) is a polynomial in the variables $\{a_{(d)}^{N}:d\in\mathbb{Z}_{\geq 1}\}$, $\{a_{\lambda}^{N}:\ell(\lambda)\geq 2\}$, but the terms that contain some $a_{\lambda}^{N}$, with $\ell(\lambda)\geq 2$, vanish after taking the limit. As a result, we can assume w.l.o.g. that $a_{\lambda}^{N}=0$, whenever $\ell(\lambda)\geq 2$, from the beginning. So, instead of $F_{N,\theta}(\vec{x})$ in (15), we can consider
$H_{N,\theta}(\vec{x}):=\sum_{d\geq 1}{a_{(d)}^{N}p_{(d)}(\vec{x})}=\sum_{d\geq 1}\sum_{i=1}^{N}{a_{(d)}^{N}x_{i}^{d}},$
and finish by proving that
$\lim_{N\to\infty}{N^{-|\mu|}\Big[\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{1}^{\mu_{1}}\Big]e^{H_{N,\theta}(\vec{x})}\,\Big|_{\vec{x}=(0^{N})}}=\prod_{i=1}^{s}{m_{\mu_{i}}}.$ (26)
At this point, the proof of (26) follows exactly from the same argument as the proof of [BCG22, Proposition 5.6] (which studied the high temperature regime $N\to\infty$, $N\theta\to\text{const}$) with only minor modifications. The conclusion from this argument is that
$N^{-|\mu|}\Big[\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{1}^{\mu_{1}}\Big]e^{H_{N,\theta}(\vec{x})}\,\Big|_{\vec{x}=(0^{N})}=\prod_{i=1}^{s}{\Big([x^{0}]\big(\theta\cdot Z+*_{h_{N}}\big)^{\mu_{i}-1}\big(h_{N}(x)\big)\Big)}+O\bigg(\frac{1}{N}\bigg),$ (27)
where $h_{N}(x):=\frac{1}{N}\cdot\partial_{1}H_{N,\theta}(\vec{x})|_{x_{1}=x}=\sum_{d=1}^{\infty}{\frac{da^{N}_{(d)}}{N}x^{d-1}}$, the operator $Z$ acts on the ring of formal power series in $x$ by $Z(x^{m}):=\mathbf{1}_{\{m=0\}}\cdot x^{m-1}$, and $_{h_{N}}$ is multiplication by $h_{N}(x)$. From (27) and the assumption (16), we have
$\lim_{N\to\infty}{N^{-|\mu|}\Big[\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{1}^{\mu_{1}}\Big]e^{H_{N,\theta}(\vec{x})}\,\Big|_{\vec{x}=(0^{N})}}=\prod_{i=1}^{s}{\Big([x^{0}]\big(\theta\cdot Z+*_{h}\big)^{\mu_{i}-1}\big(h(x)\big)\Big)},$ (28)
where $h(x):=\sum_{d=1}^{\infty}{\theta^{1-d}\kappa_{d}x^{d-1}}$.
Step 4. In view of (28), our desired equation (26) would follow from the equalities
$x^{0}]\big(\theta\cdot Z+*_{h}\big)^{\ell-1}\big(h(x)\big)=m_{\ell},\text{ for all }\ell\in\mathbb{Z}_{\geq 1},$ (29)
where $h(x):=\sum_{d=1}^{\infty}{\theta^{1-d}\kappa_{d}x^{d-1}}$ and $m_{\ell}$ is the function of $\kappa_{1},\dots,\kappa_{\ell}$ given by (18). This equality can be established by known arguments in the literature, e.g. see the proof of [Cue26, Theorem 3.6].
The argument is the following. Fix any $\ell\in\mathbb{Z}_{\geq 1}$, and express the left hand side of (29) as the sum
$x^{0}]\big(\theta\cdot Z+*_{h}\big)^{\ell-1}\big(h(x)\big)=\sum_{w_{1},w_{2},\dots,w_{\ell}}{w_{\ell}\cdots w_{2}w_{1}(1)}$ (30)
over sequences $w_{1},w_{2},\dots,w_{\ell}$ such that:
$\bullet$ each $w_{i}$ is the operator of multiplication by $\kappa_{d}x^{d-1}$, for some $d\in\mathbb{Z}_{\geq 1}$, or the operator $\theta\cdot Z$;
[…]